what is an equation of a line that is perpendicular to y=3/5x-2 and passed through point(5,-7)
Answer by stanbon(75887)

what is an equation of a line that is perpendicular to y=3/5x-2 and passed through point(5,-7)
------------------
The given line has slope = 3/5
----
A perpendicular line must have slope -5/3
----
Form::
y = mx + b
Substitute for x,y,m and solve for "b":
-7 = (-5/3)5 + b
-------
-21/3 = -25/3 + b
b = -46/3
--------
Ans: y = (-5/3)x - (46/3)
----------------------------
Cheers,
Stan H.
======================